# The Validation Gap: A Mechanistic Analysis of How Language Models Compute Arithmetic but Fail to Validate It

### v1

@@ D.1 Models
Details of the models used in this study are presented in Table 2. Specifically, we include information on the number of parameters, the number of layers, the size of the hidden model dimension, the number of attention heads per attention block, and the respective model licenses. All models are instruction-tuned and expect a series of special tokens (e.g., to indicate the beginning of the user prompt or the end of a turn). Therefore, we wrap all prompts in the respective chat templates of the models666huggingface.co/docs/transformers/chat_templating.. When applicable, we use the models’ default system prompts.
@@ D.3 Circuits for Arithmetic Computation
Model Attention Heads
Consistency Qwen-2.5-1.5B-Instruct L12H0, L12H2, L12H10, L13H0, L13H1, L13H10
Qwen-2.5-Math-1.5B-Instruct L13H0, L13H1
Llama-3.2-3B-Instruct L4H14, L5H3, L7H11, L8H1, L10H5, L10H18
Phi-3-Mini-4k-Instruct L10H1, L10H5, L10H14, L14H19, L16H18
Random Qwen-2.5-1.5B-Instruct L1H0, L4H9, L9H6, L10H5, L11H8, L27H9
Qwen-2.5-Math-1.5B-Instruct L4H9, L27H5
Llama-3.2-3B-Instruct L1H1, L4H19, L9H16, L10H15, L11H23, L27H12
Phi-3-Mini-4k-Instruct L0H25, L3H18, L8H25, L20H19, L23H28
@@ D.4 Patching Consistency Heads
To evaluate the influence of individual consistency heads on the models’ error detection behavior, we run models on prompts $X_{both}$, containing a consistent error at both the position of the arithmetic result and the final numeric answer, and patch the latent activations of the consistency heads listed in Table 3 with activations associated with prompts $X_{result}$, where the error appears only at the arithmetic result.
Specifically, for an attention head $h$, let $A^{h}(X)$ denote its attention matrix for a given prompt $X$. The patching operation we perform is defined as $A^{h}(X_{both})=\alpha\cdot A^{h}(X_{result})$, where $\alpha$ is a scaling factor that controls the influence of the patched activation. We set $\alpha$ to $3.1$ for Qwen-2.5-1.5B-Instruct and Llama-3.2-3B-Instruct, to $3.3$ and $3.4$ for Qwen-2.5-Math-1.5B-Instruct and Phi-3-Mini-4k-Instruct, respectively. We perform the patching over $1,000$ prompt pairs. As a control setup for this experiment, we compare the result to patching randomly selected attention heads that are not labeled as consistency heads (see the full list in Table 3).
@@ E.1 Error Detection Circuits
Figure 9 illustrates the faithfulness scores and number of edges for the soft intersection circuit $\mathcal{C}^{(\tau)}$ of Qwen-2.5-1.5B-Instruct across various threshold values $\tau$. Specifically, Figure 9(a) displays values for the circuit responsible for detecting errors at the position of the arithmetic results, while Figure 9(b) shows values for an error at the final numeric answer. The red circles indicate the circuits that offer the best balance between faithfulness and size. For detecting arithmetic errors at the position of the arithmetic result, the optimal threshold is $\tau=\frac{5}{8}$, while for detecting errors at the final numeric answer, the strict intersection at $\tau=\frac{8}{8}$ provides the best trade-off. The corresponding circuits are visualized in Figures 24 and 28, respectively.
Similar results for Qwen-2.5-Math-1.5B-Instruct are shown in Figure 10, where the optimal circuits are $\mathcal{C}_{\text{result}}^{(7/8)}$ and $\mathcal{C}_{\text{answer}}^{(8/8)}$, depicted in Figures 25 and 29. For Llama-3.2-3B-Instruct, the corresponding results are displayed in Figure 11, with optimal circuits $\mathcal{C}_{\text{result}}^{(7/8)}$ and $\mathcal{C}_{\text{answer}}^{(8/8)}$, visualized in Figures 26 and 30. Finally, Figure 12 provides the results for Phi-3-Mini-4k-Instruct, where the best soft intersection circuits are $\mathcal{C}_{\text{result}}^{(7/8)}$ and $\mathcal{C}_{\text{answer}}^{(6/8)}$, shown in Figures 27 and 31.
@@ E.2 Edge Overlap of Error Detection Circuits
Table 4: Edge Overlap. The edge overlap between the two error detection circuits in terms of their Intersection over Union (IoU) and Intersection over Minimum (IoM). The metrics quantify the proportion of edges shared by both circuits while considering the token position in which an edge appears.
@@ E.4 Consistency Heads
As mentioned in Section 4.2, we find that consistency heads play an important role in the model’s arithmetic error detection process. These heads exhibit high average attention scores when the digits of the arithmetic result either align or misalign with the final numeric answer. Figures 13(a) and 13(b) illustrate examples of attention scores for two of these heads in Qwen-2.5-1.5B-Instruct. Similar patterns are presented for Qwen-2.5-Math-1.5B-Instruct in Figures 14(a) and 14(b), for Llama-3.2-3B-Instruct in Figures 15(a) and 15(b), and for Phi-3-Mini-4k-Instruct in Figures 16(a) and 16(b). A comprehensive list of all identified consistency heads for each model is provided in Table 3.
@@ E.6 Computation Circuits
As outlined in Section D.3, we identify the circuits responsible for predicting the correct arithmetic result of the equations in the prompts described in Section 3. Table 9 presents the size and faithfulness of the respective circuits for all models and templates. Results for the faithfulness scores and sizes of the soft intersection circuits $\mathcal{C}^{(\tau)}$ for different threshold values $\tau$ for all models are shown in Figures 9(c), 10(c), 11(c), and 12(c), respectively. The final circuits are visualized in Figures 20, 21, 22, and 23.
@@ Appendix F Implementation Details
For the majority of our circuit identification experiments, we used the AutoCircuit library developed by Miller et al. (2024). For the remaining experiments, we relied on the TransformerLens library by Nanda and Bloom (2022). All models were loaded with bfloat16 precision. The experiments were conducted on a single A100 GPU with 80GB of memory, consuming approximately 350 GPU hours in total. Additionally, GitHub Copilot was used as an assistant tool for parts of the project’s source code development.
@@
Table 7: Problem templates. The 8 problem templates including [instruction], [person], [object], [pronoun], [num1], [num2], [num3] as variable components. While samples within a template contain the same number of tokens, samples across templates vary in length due to differences in their non-variable parts.
@@
Table 8: Template variables. The full set of variables and possible values that are used to create the data. [instruction] is the only variable that contains [correct_pair] as another variable part.
@@
Qwen-2.5-1.5B-Instruct 1 Inv. Result 84715 281 100.00
@@
Qwen-2.5-Math-1.5B-Instruct 1 Inv. Result 84715 141 99.46
@@
Llama-3.2-3B-Instruct 1 Inv. Result 389971 190 99.28
@@
Phi-3-Mini-4k-Instruct 1 Inv Result 1592881 285 99.25
@@
Table 9: Comparison of circuit faithfulness and size. The faithfulness score and the number of edges of the circuit identified for each model, template, and task. To better compare circuit sizes, we also present the total number of edges per token position for each model. Left: Qwen models (Qwen-2.5 and Qwen-2.5-Math) from templates 1–8; Right: Llama-3.2 and Phi-3-Mini models from templates 1–8.
@@
Figure 17: Additional consistency head patching results. Accuracy of models on (clean, corrupt) prompt pairs, where the clean prompt contains a consistent error at both error positions. The blue intervention bar represents the result after patching a set of consistency heads (for a list of heads, please refer to Table 3 in the Appendix). In contrast, the orange bar indicates the accuracy after patching a set of randomly chosen attention heads that are not labeled as consistency heads.
@@
Figure 18: Additional probing results. The linear probe’s accuracy across all layers of Qwen-2.5-Math-1.5B-Instruct (Figure 18(a)), Llama-3.2-3B-Instruct (Figure 18(b)), and Phi-3-Mini-4k-Instruct (Figure 18(c)) at selected token positions. Only in higher layers, the probe is able to achieve predict the correct arithmetic result perfectly. For Qwen-2.5-Math-1.5B-Instruct and Phi-3-Mini-4k-Instruct, we observe moderate accuracies also in middle layers.
@@
Figure 19: Additional results for interventions on the residual stream. Accuracy of models on (clean, corrupt) prompt pairs where (left) the clean prompt contains a consistent error at both error positions, and (right) an error is present only at the position of the arithmetic result. The blue intervention bar denotes the result after adding the hidden representation of the residual stream in layer higher layers (at [result-first]) to the residual stream in lower layers (at [result-second]). For Qwen-2.5-Math-1.5B-Instruct (Figure 19(a)), the residual steams’ hidden representation from layer 22 is added to the one in layer 1. For Llama-3.2-3B-Instruct (Figure 19(b)), we add the representation from layer 16 to layer 2, and for Phi-3-Mini-4k-Instruct (Figure 19(c)), the representation from layer 24 is added to layer 1.

### v2

@@ D.1 Models
Details of the models used in this study are presented in Table 3. Specifically, we include information on the number of parameters, the number of layers, the size of the hidden model dimension, the number of attention heads per attention block, and the respective model licenses. All models are instruction-tuned and expect a series of special tokens (e.g., to indicate the beginning of the user prompt or the end of a turn). Therefore, we wrap all prompts in the respective chat templates of the models777huggingface.co/docs/transformers/chat_templating.. When applicable, we use the models’ default system prompts.
@@ D.4 Patching Consistency Heads
To evaluate the influence of individual consistency heads on the models’ error detection behavior, we conduct two complementary patching interventions. In the first intervention, we run models on prompts $X_{both}$, which contain a consistent error at both the arithmetic result and the final numeric answer (for which models tend to incorrectly predict “valid”), and patch the latent activations of the consistency heads listed in Table 4 with activations from prompts $X_{result}$, where the error appears only at the arithmetic result (for which models typically predict “invalid”). This intervention is expected to increase the rate of “invalid” predictions. In the second, reverse intervention, we run models on prompts $X_{result}$ and patch the same heads with activations from $X_{both}$. This setup is expected to decrease the rate of “invalid” predictions.
Specifically, for an attention head $h$, let $A^{h}(X)$ denote its attention matrix for a given prompt $X$. The patching operation we perform is defined as $A^{h}(X_{target})=\alpha\cdot A^{h}(X_{source})$, where $\alpha$ is a scaling factor that controls the influence of the patched activation. For the first intervention, we set $\alpha$ to $3.1$ for Qwen-2.5-1.5B-Instruct and Llama-3.2-3B-Instruct, and to $3.3$ and $3.4$ for Qwen-2.5-Math-1.5B-Instruct and Phi-3-Mini-4k-Instruct, respectively. For the second intervention, we set $\alpha=1.0$ for all models. We perform the patching over $1,000$ prompt pairs. As a control setup for this experiment, we compare the result to patching randomly selected attention heads that are not labeled as consistency heads (see the full list in Table 4).
@@ E.1 Error Detection Circuits
Figure 10 illustrates the faithfulness scores and number of edges for the soft intersection circuit $\mathcal{C}^{(\tau)}$ of Qwen-2.5-1.5B-Instruct across various threshold values $\tau$. Specifically, Figure 10(a) displays values for the circuit responsible for detecting errors at the position of the arithmetic results, while Figure 10(b) shows values for an error at the final numeric answer. The red circles indicate the circuits that offer the best balance between faithfulness and size. For detecting arithmetic errors at the position of the arithmetic result, the optimal threshold is $\tau=\frac{5}{8}$, while for detecting errors at the final numeric answer, the strict intersection at $\tau=\frac{8}{8}$ provides the best trade-off. The corresponding circuits are visualized in Figures 26 and 30, respectively.
Similar results for Qwen-2.5-Math-1.5B-Instruct are shown in Figure 11, where the optimal circuits are $\mathcal{C}_{\text{result}}^{(7/8)}$ and $\mathcal{C}_{\text{answer}}^{(8/8)}$, depicted in Figures 27 and 31. For Llama-3.2-3B-Instruct, the corresponding results are displayed in Figure 12, with optimal circuits $\mathcal{C}_{\text{result}}^{(7/8)}$ and $\mathcal{C}_{\text{answer}}^{(8/8)}$, visualized in Figures 28 and 32. Finally, Figure 13 provides the results for Phi-3-Mini-4k-Instruct, where the best soft intersection circuits are $\mathcal{C}_{\text{result}}^{(7/8)}$ and $\mathcal{C}_{\text{answer}}^{(6/8)}$, shown in Figures 29 and 33.
@@ E.2 Edge Overlap of Error Detection Circuits
Table 5: The edge overlap between the two error detection circuits in terms of their Intersection over Union (IoU) and Intersection over Minimum (IoM). The metrics quantify the proportion of edges shared by both circuits while considering the token position in which an edge appears.
@@ E.3 Detection of Consistent Errors
To study how these findings transfer to larger models, we additionally evaluate the performance of LLaMA-3.1-70B-Instruct and Qwen-2.5-32B-Instruct on different types of errors. As shown in Table 8, we observe that, similar to the smaller models, both larger LLMs struggle more with detecting a consistent error at both error positions (result & answer) than with detecting an error present at only the arithmetic result or the final numeric answer. In particular, the average accuracy of LLaMA-3.1-70B-Instruct for consistent errors is $44.67\%\pm 10.50\%$, a drop of more than half compared to its $100.0\%\pm 0.00\%$ accuracy on detecting errors present only at the position of the arithmetic result or the final numeric answer.
Operation Error Type Qwen2.5-1.5B Llama-3.2-3B Phi-3-Mini-3.8B Qwen2.5-Math-1.5B
Subtraction Arithmetic Result 71.51 $\pm$ 11.50 99.66 $\pm$ 0.57 97.56 $\pm$ 5.74 93.54 $\pm$ 11.09
Numeric Answer 70.05 $\pm$ 11.57 99.60 $\pm$ 0.65 97.60 $\pm$ 5.68 93.61 $\pm$ 10.77
Result & Answer 8.68 $\pm$ 4.98 14.08 $\pm$ 2.01 99.64 $\pm$ 0.78 14.33 $\pm$ 1.60
Multiplication Arithmetic Result 55.30 $\pm$ 19.46 78.53 $\pm$ 31.47 99.54 $\pm$ 0.73 94.00 $\pm$ 13.68
Numeric Answer 55.22 $\pm$ 18.50 78.89 $\pm$ 31.11 99.54 $\pm$ 0.62 95.03 $\pm$ 13.12
Result & Answer 62.88 $\pm$ 17.26 42.73 $\pm$ 30.73 99.65 $\pm$ 0.45 80.96 $\pm$ 9.25
Division Arithmetic Result 59.86 $\pm$ 19.08 86.44 $\pm$ 13.04 94.52 $\pm$ 12.35 99.30 $\pm$ 0.92
Numeric Answer 60.06 $\pm$ 18.07 86.76 $\pm$ 12.97 94.18 $\pm$ 12.72 99.53 $\pm$ 0.69
Result & Answer 50.90 $\pm$ 15.84 11.79 $\pm$ 7.99 99.14 $\pm$ 1.50 57.95 $\pm$ 13.89
Table 7: Accuracy of models in correctly classifying the solutions’ validity of (clean, corrupt) prompt pairs across different arithmetic operations and error types. Values represent the mean accuracy across all templates, reported with their corresponding standard deviation. Cells highlighted in red indicate cases where the accuracy for Result & Answer is notably lower than for Numeric Answer and Arithmetic Result.
@@ E.4 Consistency Heads
As mentioned in Section 4.2, we find that consistency heads play an important role in the model’s arithmetic error detection process. These heads exhibit high average attention scores when the digits of the arithmetic result either align or misalign with the final numeric answer. Figures 14(a) and 14(b) illustrate examples of attention scores for two of these heads in Qwen-2.5-1.5B-Instruct. Similar patterns are presented for Qwen-2.5-Math-1.5B-Instruct in Figures 15(a) and 15(b), for Llama-3.2-3B-Instruct in Figures 16(a) and 16(b), and for Phi-3-Mini-4k-Instruct in Figures 17(a) and 17(b). A comprehensive list of all identified consistency heads for each model is provided in Table 4.
@@ E.6 Computation Circuits
As outlined in Section D.3, we identify the circuits responsible for predicting the correct arithmetic result of the equations in the prompts described in Section 3. Table 12 presents the size and faithfulness of the respective circuits for all models and templates. Results for the faithfulness scores and sizes of the soft intersection circuits $\mathcal{C}^{(\tau)}$ for different threshold values $\tau$ for all models are shown in Figures 10(c), 11(c), 12(c), and 13(c), respectively. The final circuits are visualized in Figures 22, 23, 24, and 25.
Error Type Qwen-2.5-32B Llama-3.1-70B
Arithmetic Result 98.93 $\pm$ 1.57 100.00 $\pm$ 0.00
Numeric Answer 99.51 $\pm$ 0.09 100.00 $\pm$ 0.00
Result & Answer 80.13 $\pm$ 9.65 44.76 $\pm$ 10.50
Table 8: Accuracy of bigger models in correctly classifying the solutions’ validity of (clean, corrupt) prompt pairs with different error types. Values represent the mean accuracy across all templates, reported with their corresponding standard deviation.
@@ Appendix F Implementation Details
For the majority of our circuit identification experiments, we used the AutoCircuit library developed by Miller et al. (2024). For the remaining experiments, we relied on the TransformerLens library by Nanda and Bloom (2022). All models were loaded with bfloat16 precision. The experiments were conducted on a single A100 GPU with 80GB of memory, consuming approximately 350 GPU hours in total. Additionally, GitHub Copilot was used as an assistant tool for parts of the project’s source code development, and ChatGPT was used to correct minor grammatical errors.
@@
Table 9: The 8 problem templates including [instruction], [person], [object], [pronoun], [num1], [num2], [num3] as variable components. While samples within a template contain the same number of tokens, samples across templates vary in length due to differences in their non-variable parts.
Subtraction Templates 1-8 Multiplication Templates 1-8 Division Templates 1-8
[instruction] Problem: [person] has [num1] [object]. [pronoun] [verb] [num2] [object]. How many [object] does [pronoun] have now? Reasoning: [person] has [num1] - [num2] = [num3] [object]. So, [pronoun] has [num3] [object] remaining. Answer: The above reasoning is [instruction] Problem: [person] has [num1] [object] per day. After [num2] days, how many [object] does [pronoun] have in total? Reasoning: [person] has [num1] × [num2] = [num3] [object]. So, [pronoun] has [num3] [object] in total. Answer: The above reasoning is [instruction] Problem: [person] has [num1] [object]. [pronoun] wants to organize them into equal groups of [num2] [object] each. How many groups can [pronoun] make? Reasoning: [person] can make [num1] ÷ [num2] = [num3] groups. So, [pronoun] can make [num3] groups. Answer: The above reasoning is
[instruction] Problem: [person] starts with [num1] [object]. After [pronoun] [verb] [num2], how many [object] does [pronoun] have left? Reasoning: To solve this, we subtract [num2] from [num1]: [num1] - [num2] = [num3]. Therefore, [person] now has [num3] [object]. Answer: The above reasoning is [instruction] Problem: [person] buys [num1] [object] each time [pronoun] goes shopping. If [pronoun] goes shopping [num2] times, how many [object] does [pronoun] buy in total? Reasoning: To solve this, we multiply [num1] and [num2]: [num1] × [num2] = [num3]. Therefore, [person] buys [num3] [object] in total. Answer: The above reasoning is [instruction] Problem: [person] starts with [num1] [object]. If [pronoun] puts [num2] [object] in each container, how many containers can [pronoun] fill? Reasoning: To solve this, we divide [num1] by [num2]: [num1] ÷ [num2] = [num3]. Therefore, [person] can fill [num3] containers. Answer: The above reasoning is
[instruction] Problem: Initially, [person] possesses [num1] [object]. [pronoun] then [verb] [num2] [object]. What’s the remaining amount of [object] that [pronoun] has? Reasoning: We calculate: [num1] (original) - [num2] (removed) = [num3] (remaining). So, [person] now has [num3] [object]. Answer: The above reasoning is [instruction] Problem: [person] collects [num1] [object] each week. After [num2] weeks, how many [object] has [pronoun] collected altogether? Reasoning: We calculate: [num1] (per week) × [num2] (weeks) = [num3] (total). So, [pronoun] has collected [num3] [object] altogether. Answer: The above reasoning is [instruction] Problem: [person] has [num1] [object] to share equally. If each person gets [num2] [object], how many people can receive [object]? Reasoning: We calculate: [num1] (total) ÷ [num2] (per person) = [num3] (people). So, [num3] people can receive [object]. Answer: The above reasoning is
[instruction] Problem: [person]’s collection of [object] decreases from [num1] to an unknown amount after [pronoun] [verb] [num2] [object]. Reasoning: To find the remaining total, we subtract: [num1] - [num2] = [num3] (final amount). Thus, [person] ends up with [num3] [object]. Answer: The above reasoning is [instruction] Problem: Initially, [person] receives [num1] [object] each month. After [num2] months, what’s the total amount of [object] that [pronoun] has received? Reasoning: We calculate: [num1] (per month) × [num2] (months) = [num3] (total). So, [person] has received [num3] [object]. Answer: The above reasoning is [instruction] Problem: Initially, [person] possesses [num1] [object]. [pronoun] wants to arrange them in rows with [num2] [object] per row. What’s the total number of rows that can be formed? Reasoning: We calculate: [num1] (total) ÷ [num2] (per row) = [num3] (rows). So, [person] can form [num3] rows. Answer: The above reasoning is
[instruction] Problem: [person] originally owns [num1] [object]. After [pronoun] [verb] [num2] [object], how many does [pronoun] have left? Reasoning: A simple subtraction gives us [num1] - [num2] = [num3]. Therefore, [person] has [num3] [object] remaining. Answer: The above reasoning is [instruction] Problem: [person] originally gets [num1] [object] per visit. After [num2] visits, how many [object] has [pronoun] gotten altogether? Reasoning: A simple multiplication gives us [num1] × [num2] = [num3]. Therefore, [person] has gotten [num3] [object] in total. Answer: The above reasoning is [instruction] Problem: [person] originally owns [num1] [object]. If [pronoun] distributes [num2] [object] to each recipient, how many recipients can get [object]? Reasoning: A simple division gives us [num1] ÷ [num2] = [num3]. Therefore, [num3] recipients can get [object]. Answer: The above reasoning is
[instruction] Problem: [person] possesses [num1] [object] at first. If [pronoun] [verb] [num2] [object], what is the remaining count? Reasoning: Subtracting them gives: [num1] - [num2] = [num3]. Consequently, [person] has [num3] [object] left. Answer: The above reasoning is [instruction] Problem: [person] earns [num1] [object] per task at first. If [pronoun] completes [num2] tasks, what is the total count of [object]? Reasoning: Multiplying them gives: [num1] × [num2] = [num3]. Consequently, [pronoun] earns [num3] [object] in total. Answer: The above reasoning is [instruction] Problem: [person] possesses [num1] [object] at first. If [pronoun] arranges [num2] [object] in each pile, what is the total number of piles? Reasoning: Dividing them gives: [num1] ÷ [num2] = [num3]. Consequently, [person] can make [num3] piles. Answer: The above reasoning is
[instruction] Problem: [num1] [object] belong to [person]. [pronoun] [verb] [num2] of them. What’s the remainder? Reasoning: By subtraction, we get [num1] - [num2] = [num3]. Thus, [person] has [num3] [object] remaining. Answer: The above reasoning is [instruction] Problem: [person] finds [num1] [object] each time [pronoun] searches. After [num2] searches, what’s the total? Reasoning: By multiplication, we get [num1] × [num2] = [num3]. Thus, [person] finds [num3] [object] in total. Answer: The above reasoning is [instruction] Problem: [num1] [object] belong to [person]. [pronoun] places [num2] [object] in each box. What’s the total number of boxes needed? Reasoning: By division, we get [num1] ÷ [num2] = [num3]. Thus, [person] needs [num3] boxes. Answer: The above reasoning is
[instruction] Problem: [person] begins with [num1] [object] and then [verb] [num2] of them. How many [object] does [pronoun] have left? Reasoning: Let’s subtract them: [num1] - [num2] = [num3]. Therefore, [person] has [num3] [object] remaining. Answer: The above reasoning is [instruction] Problem: [person] produces [num1] [object] per session and then has [num2] sessions. How many [object] are there in total? Reasoning: Let’s multiply them: [num1] × [num2] = [num3]. Therefore, there are [num3] [object] altogether. Answer: The above reasoning is [instruction] Problem: [person] begins with [num1] [object] and then organizes [num2] [object] per shelf. How many shelves does [pronoun] need? Reasoning: Let’s divide them: [num1] ÷ [num2] = [num3]. Therefore, [person] needs [num3] shelves. Answer: The above reasoning is
Table 10: Templates used for other arithmetic operations. We created 8 templates for each operation, including [instruction], [person], [object], [pronoun], [num1], [num2], [num3] as variable components. The subtraction template also has a [verb] variable, for which we use the following verbs: “lost”, “sold”, “gave away”, “donated”, “threw away”,
@@
Table 11: The full set of variables and possible values that are used to create the data. [instruction] is the only variable that contains [correct_pair] as another variable part.
@@
Qwen-2.5-1.5B-Instruct
1 Inv. Result 84715 281 100.00
@@
Qwen-2.5-Math-1.5B-Instruct
1 Inv. Result 84715 141 99.46
@@
Llama-3.2-3B-Instruct
1 Inv. Result 389971 190 99.28
@@
Phi-3-Mini-4k-Instruct
1 Inv Result 1592881 285 99.25
@@
Table 12: The faithfulness score and the number of edges of the circuit identified for each model, template, and task. To better compare circuit sizes, we also present the total number of edges per token position for each model. Left: Qwen models (Qwen-2.5 and Qwen-2.5-Math) from templates 1–8; Right: Llama-3.2 and Phi-3-Mini models from templates 1–8.
@@
Figure 18: Accuracy of models on (clean, corrupt) prompt pairs, where the clean prompt contains a consistent error at both error positions. The blue intervention bar represents the result after patching a set of consistency heads (for a list of heads, please refer to Table 4 in the Appendix). In contrast, the orange bar indicates the accuracy after patching a set of randomly chosen attention heads that are not labeled as consistency heads.
(a) Qwen-2.5-Math-1.5B-Instruct
(b) Llama-3.2-3B-Instruct
Figure 19: Accuracy of models on (clean, corrupt) prompt pairs for other arithmetic operations, where the clean prompt contains a consistent error at both error positions. The blue intervention bar represents the result after patching a set of consistency heads (for a list of heads, please refer to Table 4 in the Appendix). In contrast, the red bar indicates the accuracy after patching a set of randomly chosen attention heads that are not labeled as consistency heads.
@@
Figure 20: The linear probe’s accuracy across all layers of Qwen-2.5-Math-1.5B-Instruct (Figure 20(a)), Llama-3.2-3B-Instruct (Figure 20(b)), and Phi-3-Mini-4k-Instruct (Figure 20(c)) at selected token positions. Only in higher layers, the probe is able to achieve predict the correct arithmetic result perfectly. For Qwen-2.5-Math-1.5B-Instruct and Phi-3-Mini-4k-Instruct, we observe moderate accuracies also in middle layers.
@@
Figure 21: Accuracy of models on (clean, corrupt) prompt pairs where (left) the clean prompt contains a consistent error at both error positions (invalid result & answer), and (right) an error is present only at the position of the arithmetic result (invalid result). The blue intervention bar denotes the result after adding the hidden representation of the residual stream in layer higher layers (at [result-first]) to the residual stream in lower layers (at [result-second]). For Qwen-2.5-Math-1.5B-Instruct (Figure 21(a)), the residual steams’ hidden representation from layer 22 is added to the one in layer 1. For Llama-3.2-3B-Instruct (Figure 21(b)), we add the representation from layer 16 to layer 2, and for Phi-3-Mini-4k-Instruct (Figure 21(c)), the representation from layer 24 is added to layer 1.